OpenStudy (anonymous):

what is the solution to the rational equation 4x/3 - 1/3 =5

OpenStudy (anonymous):

is it this problem: \(\large \frac{4x}{3}-\frac{1}{3}=5 \) if yes, notice the fractions already have common denominators. can you combine them as 1 fraction on the left side?

OpenStudy (anonymous):

4X-1=15 X=4

OpenStudy (anonymous):

@dpaInc yeah thats the problem so would it be (4x-1)/3 = 5 (4x-1) = 15 (if you multiply both sides by 3) then 4x= 16 (adding the -1) then divide by 4 & get x=4?

OpenStudy (ledah):

I was thinking multiply both sides by 5 but I don't think that right

OpenStudy (anonymous):

i think its 4. that seems to make sense. i have ONE more quetion similar to this one ...

OpenStudy (anonymous):

good work... and as u can see there are other ways of doing it but ultimately you'll get the same answer...:)

OpenStudy (anonymous):

4x/3-1/3=5 take LCM 4x-1 ----- =5 3 4x-1=5*3 4x-1=15 4x=15+1 4x=16 x=16/4 x=4 hope you understand :)

OpenStudy (anonymous):

thanks guys :) i understand that one :) heres the other one thats confusing: simplify 1/x + 2/9 = 1/3 i cant get it because it doesnt have a common denominator like the other one did